# Scaling the glassy dynamics of active particles: Tunable fragility and reentrance

### v1

@@ III Characterising the glassy dynamics
[…]
We show the time dependence of $F_{s}(k,t)$ and MSD($t$) at constant $\phi=0.65$ and $\tau_{p}=10^{-2}$ with varying $f_{0}^{2}$ in Figs. 1(a, b). The self-intermediate scattering function $F_{s}(k,t)$ first decays towards a non-zero plateau value and at later times to zero, see Fig. 1(a). The plateau region of $F_{s}(k,t)$ corresponds to sub-diffusive motion in the MSD, which only becomes diffusive at much longer times, see Fig. 1(b). As $f_{0}$ increases, the decay of $F_{s}(k,t)$ and the transition in the MSD from sub-diffusive to diffusive behavior occur at shorter times. Overall the dynamical characteristics of $F_{s}(k,t)$ and MSD with increasing $f_{0}$ are similar to those of equilibrium supercooled fluids with increasing temperature $T$. For further quantification one can define the relaxation time $\tau_{\alpha}$ as the value of $t$ for which $F_{s}(k,t)$ decays to $1/e$, explicitly $F_{s}(k,\tau_{\alpha})=1/e$. From Fig. 1(a) is then clear that $\tau_{\alpha}$ decreases as $f_{0}$ increases. We have explored the dynamics at several other packing fractions $\phi$, and the qualitative behavior remains the same.
@@
The results in this section are consistent with existing literature on active glassy systems, where the effects of activity remain equilibrium-like at a suitably defined effective temperature when $\tau_{p}$ is small Mandal et al. (2016); Flenner et al. (2016); Nandi and Gov (2017); Nandi et al. (2018). However, we will show that this scenario drastically changes when $\tau_{p}$ is varied over a broader range and that nontrivial effects are revealed when varying $\tau_{p}$.
@@ IV Three-dimensional phase diagram
We now construct the three-dimensional liquid-glass phase diagram in the full control parameter space $(f_{0},\tau_{p},\phi)$ to understand how the persistence time $\tau_{p}$ changes the results set out in Sec. III above. For this purpose, we obtain the critical lines $\phi_{c}(f_{0}^{2},\tau_{p})$ by repeating the fitting procedure shown in Fig. 1(c) for several values of $\tau_{p}$. This numerical exploration of the three-dimensional space represents a significant computational effort, given that it requires simulations scanning a range of volume fractions $\phi$ for each pair of values $(f_{0},\tau_{p})$.
The symbols in Fig. 2(a) show the resulting critical lines, $\phi_{c}(f_{0}^{2},\tau_{p})$ in the $(f_{0},\phi)$ plane for several values of $\tau_{p}$. Similar data were obtained in Ref. Berthier et al. (2017) using the diffusivity, but the detailed characteristics and consequences were not fully explored.
@@
Qualitatively this behavior arises because of the opposite trends mentioned above: the critical liquid-glass density increases with $\tau_{p}$ at small $f_{0}$, whereas it decreases at large $f_{0}$. These contrasting trends create an intermediate range of $f_{0}$ where the critical density is non-monotonic in $\tau_{p}$. We show this explicitly in Figs. 3(b, c, d) where we follow the evolution with $\tau_{p}$ of $\phi_{c}(f_{0}^{2},\tau_{p})$ for a small, intermediate and large value of $f_{0}$ and observe increasing, non-monotonic, and decreasing variation with $\tau_{p}$, respectively. ‘
@@
This critical density is, by construction, controlled by the second term when $f_{0}\to 0$, so that $\phi_{c}$ increases with $\tau_{p}$ along with $\phi_{d}$. For large $f_{0}$, on the other hand, the first term dominates and $\phi_{c}$ then decreases with $\tau_{p}$ due to the increase of $a$ (recall Fig. S3 in SM). The lines in Figs. 3(b, c, d) confirm that the empirical description of the glass surface obtained in Sec. IV describe the numerical data, and thus the re-entrant dynamics, very well. Using the VFT-like relation between $\tau_{\alpha}$ and $\phi_{c}$ in Eq. (6), the non-monotonic behaviour of $\phi_{c}$ directly accounts for the re-entrant dynamics, as demonstrated by the full lines in Fig. 3(a) [see Sec. S8 in the SM for more details].
@@
The re-entrant dynamics observed when $\tau_{p}$ increases is accompanied by a qualitative change in the physical relaxation process, since the system crosses over from near equilibrium glassy relaxation dynamics when $\tau_{p}\to 0$ to non-thermal jamming physics when $\tau_{p}\to\infty$, as explained in Sec. II. This evolution from glass to jamming physics is captured by the behaviour of the self-intermediate scattering function shown in Fig. 3(e), which we show here for constant $\phi=0.65$ and $f_{0}^{2}=0.045$ while varying $\tau_{p}$ over a broad range of five orders of magnitude. At small $\tau_{p}$, $F_{s}(k,t)$ shows the characteristic two-step relaxation decay typical of glassy relaxation. However, at very large $\tau_{p}$, $F_{s}(k,t)$ decays in a single step with no intermediate plateau, as usually observed in driven systems in the absence of thermal fluctuations Ikeda et al. (2012, 2013). As expected, the MSD also shows a similar change of behavior with changing $\tau_{p}$ (SM Fig. S6). Beyond these qualitative changes, the non-monotonic dependence on $\tau_{p}$ of the final decay time of $F_{s}(k,t)$ is also evident from Fig. 3(e).
@@ VI Tunable fragility and sub-Arrhenius to super-Arrhenius crossover
A qualitatively similar evolution with the density of the glass fragility was observed previously in the Brownian limit, $\tau_{p}\to 0$, for a similar model of soft repulsive spheres Berthier and Witten (2009a, b). The qualitative explanation is relatively straightforward. When decreasing $T$ at constant $\phi$, the system ends up in an equilibrium hard sphere fluid if $\phi<\phi_{d}$. As a result, $\tau_{\alpha}(T\to 0)$ does not diverge, and this qualitatively explains the apparent sub-Arrhenius behaviour. Instead, above the critical density $\phi_{d}$, the system crosses a glass transition at a finite $T$, and as a consequence the relaxation dynamics appears to diverge at a finite $T$, which gives rise to a super-Arrhenius temperature dependence. In this view, the glass fragility is changing continuously with density, and it is primarily controlled by the distance to the critical density $\phi_{d}$ that characterises the $T\to 0$ hard sphere limit.
@@ VII Dynamic scaling near the hard sphere non-equilibrium glass transition
Figure 5: Dynamic scaling analysis collapses the glassy dynamics of active particles. (a) Angell plot using the rescaled realxation time $\tau_{\alpha}\sqrt{T_{\text{eff}}}$ as a function of $1/T_{\text{eff}}$. Different symbols are for different $\phi$, different colors are for different $\tau_{p}$. (b) Global data collapse along the two branches describing the $\phi<\phi_{0}$ sub-Arrhenius and $\phi>\phi_{0}$ super-Arrhenius family of curves, as described in Eq. (11). (c) The exponents $\delta$ and $\mu$ depart weakly from their equilibrium value as $\tau_{p}$ increases. (d) The critical hard sphere density $\phi_{0}(\tau_{p})$ changes smoothly with $\tau_{p}$.
@@
We are now in a position to apply Eq. (11) to our data. In practice, we find that the largest value of the packing fraction, $\phi=0.8$, is too far above the critical density $\phi_{0}$ and does not provide a good data collapse. Therefore, we did not use these data for the scaling analysis. The data collapse procedure is somewhat tedious as it requires the simultaneous identification of the free parameters $\delta$, $\mu$ and $\phi_{0}$ for each $\tau_{p}$. In practice, we initialized the fitting process with the values obtained in equilibrium Berthier and Witten (2009b) and slowly varied the fit parameters to achieve a satisfying result for all $\tau_{p}$ values. The outcome of this analysis is shown in Fig. 5(b), showing the data collapse of the relaxation data along the two branches described by $F_{-}$ and $F_{+}$, while the fitted parameters are shown in Figs. 5(c, d). The exponents $\delta$ and $\mu$ vary weakly with $\tau_{p}$ and show only small departures from their equilibrium values (corresponding to $\tau_{p}\to 0$). The critical density $\phi_{0}$ increases gradually with $\tau_{p}$ and saturates at large persistence times. As expected, it closely mirrors the evolution of $\phi_{d}$ discussed above (see SM Fig. S4), as they both describe the same hard sphere dynamics using slightly different functional forms, Eqs. (6, 10).
@@ VIII Discussion and perspectives
We have studied the glassy dynamics in an active system of self-propelled soft spherical particles. This model system interpolates between particulate equilibrium glasses with soft and hard pair interactions, and confluent biological tissues. Furthermore, the system is interesting in itself for its rich phenomenology, which we reveal here by performing a full exploration of the three-dimensional phase diagram $(f_{0},\tau_{p},\phi)$.
@@
The construction and quantitative analysis of the three-dimensional phase diagram and its various limits allowed us to account for two non-trivial dynamic features: re-entrant glassy dynamics that emerges when $\tau_{p}$ is varied at fixed $\phi$ and $f_{0}$, and a glass fragility that is tuned by both changing $\tau_{p}$ and $\phi$. Our analysis generalises, and provides a simple interpretation for, related previous reports of anomalous dynamics in self-propelled particle systems Berthier et al. (2017); Arora et al. (2022); Debets et al. (2021). Our central conclusion is that the effect of activity is very well captured by the known equilibrium scaling description Berthier and Witten (2009b), provided one promotes the hard sphere glass point $\phi_{0}$ to a hard sphere glass line $\phi_{0}(\tau_{p})$ that depends continuously on the persistence time. Interestingly, along this glass line of the active system one is effectively moving smoothly from a glass transition for small persistence times $\tau_{p}$ to a jamming transition for large $\tau_{p}$ Mandal et al. (2020); Mandal and Sollich (2020); Wiese et al. (2023) as shown by the disappearance of the intermediate plateau in the relaxation functions with increasing $\tau_{p}$.
@@
An interesting perspective for future research is to connect the behaviour of soft active particles we have studied here to the physics revealed by studies of model systems for confluent epithelial tissues. In these models, the packing fraction $\phi$ does not control the physics as in soft particles as it is effectively fixed to unity. Instead there is a structural parameter controlling the behaviour of the system, the so-called target perimeter $p_{0}$ Bi et al. (2016); Sadhukhan and Nandi (2021); Sadhukhan et al. (2024a); Atia et al. (2021). This drives the system from fluid-like states at large $p_{0}$ to solid-like behaviour at low $p_{0}$, in a manner analogous to $1/\phi$ in particle systems. Intriguingly, there are several reports of sub-Arrhenius to super-Arrhenius crossover in the literature as $p_{0}$ is varied Sussman et al. (2018); Sadhukhan and Nandi (2021); Li et al. (2021); Sadhukhan et al. (2024b), suggesting a possible analogy with soft particles. re-entrant dynamics also exists in variants of these model systems Arora et al. (2024). Future work should explore whether the analogy can be made more quantitative, and whether the analysis carried out here can also be useful to rationalise the characteristics of the glassy dynamics of biological tissues and its interplay with jamming physics.
Acknowledgements.

### v2

@@ III Characterising the glassy dynamics
where $N$ is the total number of particles, $\mathbf{r}_{i}(t)$ is the position of the $i$th particle at time $t$, and $\langle\ldots\rangle$ denotes an average over independent configurations and over the time origins $t_{0}$ taken at steady state. For convenience we choose $k$ as the position of the first peak of the static structure factor of the passive system at $\phi=0.65$; this first peak position does not change significantly at other values of $\phi$ [SM Fig. S1]. We first ask how the dynamics varies with $f_{0}$ at constant $\tau_{p}$, and then explore the evolution with $\tau_{p}$.
We show the time dependence of $F_{s}(k,t)$ and MSD($t$) at constant $\phi=0.65$ and $\tau_{p}=10^{-2}$ with varying $f_{0}^{2}$ in Figs. 1(a, b). The self-intermediate scattering function $F_{s}(k,t)$ first decays towards a non-zero plateau value and at later times to zero, see Fig. 1(a). The plateau region of $F_{s}(k,t)$ corresponds to sub-diffusive motion in the MSD, which only becomes diffusive at much longer times, see Fig. 1(b). As $f_{0}$ increases, the decay of $F_{s}(k,t)$ and the transition in the MSD from sub-diffusive to diffusive behavior occur at shorter times. Overall the dynamical characteristics of $F_{s}(k,t)$ and MSD with increasing $f_{0}$ are similar to those of equilibrium supercooled fluids with increasing temperature $T$. For further quantification one can define the relaxation time $\tau_{\alpha}$ as the value of $t$ for which $F_{s}(k,t)$ decays to $1/e$, explicitly $F_{s}(k,\tau_{\alpha})=1/e$. From Fig. 1(a), it is then clear that $\tau_{\alpha}$ decreases as $f_{0}$ increases. We have explored the dynamics at several other packing fractions $\phi$, and the qualitative behavior remains the same.
@@
The results in this section are broadly consistent with existing literature on active glassy systems, where the effects of activity remain equilibrium-like at a suitably defined effective temperature as long as $\tau_{p}$ is small Mandal et al. (2016); Flenner et al. (2016); Nandi and Gov (2017); Nandi et al. (2018). In the next sections, we explore a much broader range of control parameters, which will allow us to reveal additional effects that are non-trivial.
@@ IV Three-dimensional phase diagram
We now construct the three-dimensional liquid-glass phase diagram in the full control parameter space $(f_{0},\tau_{p},\phi)$ to understand how the persistence time $\tau_{p}$ changes the results set out in Sec. III above. For this purpose, we obtain the critical lines $\phi_{c}(f_{0}^{2},\tau_{p})$ by repeating the fitting procedure shown in Fig. 1(c) for several values of $\tau_{p}$. This numerical exploration of the three-dimensional space represents a significant computational effort, given that it requires simulations scanning a range of volume fractions $\phi$ for each pair of values $(f_{0},\tau_{p})$. This large computational effort is useful, as it allows us to recover within a single approach several results obtained independently in various limits, which we can rationalise and organise in the scaling approach developed below. The symbols in Fig. 2(a) show the resulting critical lines, $\phi_{c}(f_{0}^{2},\tau_{p})$ in the $(f_{0},\phi)$ plane for several values of $\tau_{p}$. Similar data were obtained in Ref. Berthier et al. (2017) using the diffusivity, but the detailed characteristics and consequences were not fully explored.
@@ V Re-entrant relaxation dynamics
Qualitatively this behavior arises because of the opposite trends mentioned above: the critical liquid-glass density increases with $\tau_{p}$ at small $f_{0}$, whereas it decreases at large $f_{0}$. These contrasting trends create an intermediate range of $f_{0}$ where the critical density is non-monotonic in $\tau_{p}$. We show this explicitly in Figs. 3(b, c, d) where we follow the evolution with $\tau_{p}$ of $\phi_{c}(f_{0}^{2},\tau_{p})$ for a small, intermediate and large value of $f_{0}$ and observe increasing, non-monotonic, and decreasing variation with $\tau_{p}$, respectively.
@@
This critical density is, by construction, controlled by the second term when $f_{0}\to 0$, so that $\phi_{c}$ increases with $\tau_{p}$ along with $\phi_{d}$. For large $f_{0}$, on the other hand, the first term dominates and $\phi_{c}$ then decreases with $\tau_{p}$ due to the increase of $a$ (recall Fig. S3 in SM). The lines in Figs. 3(b, c, d) confirm that the empirical description of the glass surface obtained in Sec. IV describes the numerical data, and thus the re-entrant dynamics, very well.
To conclusively establish that the shape of the critical surface does describe the re-entrant behavior, we use Eq. (6). This expression can be obtained within the random-first order transition theory, which constitutes a microscopic theory of the glass transition Lubchenko and Wolynes (2007); Kirkpatrick and Thirumalai (2015) that has been extended for active systems Nandi et al. (2018); Mandal et al. (2022). Using $\phi_{c}(f_{0}^{2},\tau_{p})$ in Eq. (6), we obtain $\tau_{\alpha}$ as
$\ln\tau_{\alpha}=A+\frac{B}{c[f_{0}^{2}/a(\tau_{p})]^{\beta(\tau_{p})/2}+\phi_{d}(\tau_{p})-\phi},$ (10)
where we have included the fitting parameter $c$ to account for the activity-dependence of $A$ and $B$ (the value of $c$ is nearly 1, see SM Sec. S8 for further details). The lines in Fig. 3(a) are the plots of Eq. (10) with the values of $c$ as given in Table S1. The excellent agreement between the simulation results and Eq. (10) demonstrates that it is indeed the critical surface shape that governs the re-entrant behavior. Indeed, neglecting the fit factor $c$ (see SM, Sec. S8), the key first two terms in the denominator of Eq. [10] just correspond to $\phi_{c}$. The competition between these terms determines, for given $f_{0}^{2}$ and $\phi$, the location $\tau_{p,\rm min}$ of the re-entrant minimum in $\tau_{\alpha}$ (see Fig. 3a). To understand qualitatively how $\tau_{p,\rm min}$ varies with $f_{0}$, notice from Fig. 3(b) that, for small $f_{0}$, $\phi_{c}$ is increasing with $\tau_{p}$ and hence $\tau_{\alpha}$ is decreasing, corresponding to $\tau_{p,\rm min}\to\infty$. For large $f_{0}$ the situation is reversed (Fig. 3d) and $\tau_{p,\rm min}\to 0$. In the intermediate range of $f_{0}$ where $\tau_{p,\rm min}$ is finite, it must therefore decrease with increasing $f_{0}$, consistent with the trend visible in Fig. 3(a).
@@
The re-entrant dynamics observed when $\tau_{p}$ increases is accompanied by a qualitative change in the physical relaxation process, since the system crosses over from near-equilibrium glassy relaxation dynamics when $\tau_{p}\to 0$ to non-thermal jamming physics when $\tau_{p}\to\infty$, as explained in Sec. II. This evolution from glass to jamming physics is captured by the behaviour of the self-intermediate scattering function shown in Fig. 3(e), which we show here for constant $\phi=0.65$ and $f_{0}^{2}=0.045$ while varying $\tau_{p}$ over a broad range of five orders of magnitude. At small $\tau_{p}$, $F_{s}(k,t)$ shows the characteristic two-step relaxation decay typical of glassy relaxation. However, at very large $\tau_{p}$, $F_{s}(k,t)$ decays in a single step with no intermediate plateau, as usually observed in driven systems in the absence of thermal fluctuations Ikeda et al. (2012, 2013). As expected, the MSD also shows a similar change of behavior with changing $\tau_{p}$ (SM Fig. S6). Beyond these qualitative changes, the non-monotonic dependence on $\tau_{p}$ of the final decay time of $F_{s}(k,t)$ is also evident from Fig. 3(e).
Re-entrant behavior appears in a variety of systems, including colloidal suspensions of sticky hard spheres Pham et al. (2002) and fluids confined within periodic potentials Nandi et al. (2011); Mandal et al. (2014). In these equilibrium colloidal systems, re-entrance is typically governed by large changes in the static structure. By contrast, re-entrant behavior in active systems has been described via changes in the caging dynamics Liluashvili et al. (2017); Debets et al. (2021a); Klongvessa et al. (2019) or via effective attractive interactions Arora et al. (2022); Berthier et al. (2017). Our work provides a complementary perspective on the mechanism underlying re-entrant behavior in active systems.
@@ VI Tunable fragility and sub-Arrhenius to super-Arrhenius crossover
A qualitatively similar evolution with the density of the glass fragility was observed previously in the Brownian limit, $\tau_{p}\to 0$, for a similar model of soft repulsive spheres Berthier and Witten (2009a, b). The microscopic explanation is relatively straightforward. When decreasing $T$ at constant $\phi$, the system ends up in an equilibrium hard sphere fluid if $\phi<\phi_{d}$. As a result, $\tau_{\alpha}(T\to 0)$ does not diverge, and this qualitatively explains the apparent sub-Arrhenius behaviour. Instead, above the critical density $\phi_{d}$, the system crosses a glass transition at a finite $T$, and as a consequence the relaxation dynamics appears to diverge at a finite $T$, which gives rise to a super-Arrhenius temperature dependence. In this view, the glass fragility is changing continuously with density, and it is primarily controlled by the distance to the critical density $\phi_{d}$ that characterises the $T\to 0$ hard sphere limit. Note that in this physical explanation of the evolution of fragility, the softness of the particles plays no role. In Brownian colloidal systems, particle softness has likewise been shown to play only a limited role in directly controlling glass fragility Philippe et al. (2018). In some soft charged colloids, osmotic deswelling has been shown to produce a large fragility change Mattsson et al. (2009); Pelaez-Fernandez et al. (2015), but this is unrelated to our observations.
@@ VII Dynamic scaling near the hard sphere non-equilibrium glass transition
Figure 5: Dynamic scaling analysis collapses the glassy dynamics of active particles. (a) Angell plot using the rescaled relaxation time $\tau_{\alpha}\sqrt{T_{\text{eff}}}$ as a function of $1/T_{\text{eff}}$. Different symbols are for different $\phi$, different colors are for different $\tau_{p}$. (b) Global data collapse along the two branches describing the $\phi<\phi_{0}$ sub-Arrhenius and $\phi>\phi_{0}$ super-Arrhenius family of curves, as described in Eq. (12). (c) The exponents $\delta$ and $\mu$ depart weakly from their equilibrium value as $\tau_{p}$ increases. (d) The critical hard sphere density $\phi_{0}(\tau_{p})$ changes smoothly with $\tau_{p}$.
@@
We are now in a position to apply Eq. (12) to our data. In practice, we find that the largest value of the packing fraction, $\phi=0.8$, is too far above the critical density $\phi_{0}$ and does not provide a good data collapse. Therefore, we did not use these data for the scaling analysis. The data collapse procedure is somewhat tedious as it requires the simultaneous identification of the free parameters $\delta$, $\mu$ and $\phi_{0}$ for each $\tau_{p}$. In practice, we initialized the fitting process with the values obtained in equilibrium Berthier and Witten (2009b) and slowly varied the fit parameters to achieve a satisfying result for all $\tau_{p}$ values. The outcome of this analysis is shown in Fig. 5(b), showing the data collapse of the relaxation data along the two branches described by $F_{-}$ and $F_{+}$, while the fitted parameters are shown in Figs. 5(c, d). The exponents $\delta$ and $\mu$ vary weakly with $\tau_{p}$ and show only small departures from their equilibrium values (corresponding to $\tau_{p}\to 0$). The critical density $\phi_{0}$ increases gradually with $\tau_{p}$ and saturates at large persistence times. As expected, it closely mirrors the evolution of $\phi_{d}$ discussed above (see SM Fig. S4), as they both describe the same hard sphere dynamics using slightly different functional forms, Eqs. (6, 11).
@@ VIII Discussion and perspectives
We have studied the glassy dynamics in an active system of self-propelled soft spherical particles. This model system contains as limit cases thermal soft spheres and persistent self-propelled hard spheres, while for large $f_{0}$ and $\tau_{p}$ it describes self-propelled soft particles and so connects qualitatively also to the physics of confluent biological tissues. This broad range of physical behaviours captured by the model leads to a rich phenomenology, which we reveal here by performing a full exploration of the three-dimensional phase diagram $(f_{0},\tau_{p},\phi)$.
@@
The construction and quantitative analysis of the three-dimensional phase diagram and its various limits allowed us to account for two non-trivial dynamic features: re-entrant glassy dynamics that emerges when $\tau_{p}$ is varied at fixed $\phi$ and $f_{0}$, and a glass fragility that is tuned by both changing $\tau_{p}$ and $\phi$. Our analysis generalises, and provides a simple interpretation for, related previous reports of anomalous dynamics in self-propelled particle systems Berthier et al. (2017); Arora et al. (2022); Debets et al. (2021a). Our central conclusion is that the effect of activity is very well captured by the known equilibrium scaling description Berthier and Witten (2009b), provided one promotes the hard sphere glass point $\phi_{0}$ to a hard sphere glass line $\phi_{0}(\tau_{p})$ that depends continuously on the persistence time.
Interestingly, along this glass line of the active system one is effectively moving smoothly from a glass transition for small persistence times $\tau_{p}$ to a jamming transition for large $\tau_{p}$ Mandal et al. (2020); Mandal and Sollich (2020); Wiese et al. (2023); Szamel and Flenner (2024) as shown by the disappearance of the intermediate plateau in the relaxation functions with increasing $\tau_{p}$. Microscopically, this behaviour can be rationalized by analogy with the physics of passive glasses subjected to mechanical deformation by steady shear Ikeda et al. (2012, 2013): here one also finds a smooth change between two distinct regimes, controlled by temperature. When thermal fluctuations are significant, corresponding to our active glasses at small $\tau_{p}$, relaxation under sufficiently slow shear proceeds by thermal activation across energy barriers. For times shorter than the barrier crossing time, particles can only relax partially by “rattling” in cages formed by their neighbours, causing plateaus in typical relaxation functions. In the athermal regime, on the other hand, relaxation is driven by barriers disappearing via saddle-node bifurcations as the energy landscape is gradually deformed by the applied shear. For our active systems, this is analogous to the slow tilting of the energy landscape by active forces Mandal and Sollich (2020, 2021); Keta et al. (2023) in the large $\tau_{p}$ regime; either way, the relaxation has no analog of particles rattling in cages at early times, and relaxation functions therefore do not show plateaus.
@@
An interesting perspective for future research is to connect the behaviour of soft active particles we have studied here to the physics revealed by studies of model systems for confluent epithelial tissues. In these models, the packing fraction $\phi$ does not control the physics as in soft particles as it is effectively fixed to unity. Instead there is a structural parameter controlling the behaviour of the system, the so-called target perimeter $p_{0}$ Bi et al. (2016); Sadhukhan and Nandi (2021); Sadhukhan et al. (2024a); Atia et al. (2021). This drives the system from fluid-like states at large $p_{0}$ to solid-like behaviour at low $p_{0}$, in a manner analogous to $1/\phi$ in particle systems. In spite of these important differences, the parameters $p_{0}$ and $\phi$ play similar roles in so far as they control the transition from fluid to solid response, in the absence of thermal fluctuations and active forces. Intriguingly, there are several reports of sub-Arrhenius to super-Arrhenius crossover in the literature as $p_{0}$ is varied Sussman et al. (2018); Sadhukhan and Nandi (2021); Li et al. (2021); Sadhukhan et al. (2024b), suggesting a possible analogy with soft particles. Re-entrant dynamics also exists in variants of these model systems Arora et al. (2024). Future work should explore whether the analogy can be made more quantitative, and whether the analysis carried out here can also be useful to rationalise the characteristics of the glassy dynamics of biological tissues and its interplay with jamming physics, thus hopefully illuminating the role of many-body forces and confluence in the physics of tissue models.
acknowledgement